OpenStudy (anonymous):

What is the value of x? x + 2/15 = x + 1/ 5

hero (hero):

I am assuming you mean: \[\frac{x+2}{15} = \frac{x+1}{5}\]

OpenStudy (anonymous):

x= -1/2

hero (hero):

If so, then cross multply to get: 5(x+2) = 15(x+1)

hero (hero):

After that, distribute the 5 and the 15 to get 5x + 10 = 15x + 15

hero (hero):

Now put like terms on the same side: 10 - 15 = 15x - 5x

hero (hero):

Then simplify: -5 = 10x and continue solving for x: -5/10 = x -1/2 = x

OpenStudy (anonymous):

@Hero would please answer following question for me please. Determine whether each function is even, odd, or neither. a) y= x is cubed sinx b)y=x+ sin(squared)x

hero (hero):

To determine if a function is even or odd, set x = -x, then simplify If f(x) = -f(x) then the function is a reflection over the y-axis (even) If f(x) = -f(-x) then the function is a reflection over the origin (odd)

hero (hero):

Afterwards, Input some test points to make sure.

hero (hero):

By making a table of values.